### v1

# Staking out the Proton Drip-Line of Thulium at the N=82 Shell Closure

Abstract
Direct observation of proton emission with very small emission energy is often unfeasible due to the long partial half-lives associated with tunneling through the Coulomb barrier. Therefore proton emitters with very small Q-values may require masses of both parent and daughter nuclei to establish them as proton unbound. Nuclear mass models have been used to predict the proton drip-line of the thulium (Tm) isotopic chain ($Z=69$), but up until now the proton separation energy has not been experimentally tested. Mass measurements were therefore performed using a Multiple Reflection Time-Of-Flight Mass Spectrometer [1] (MR-TOF-MS) at TRIUMF’s TITAN facility to definitively map the limit of proton-bound Tm. The masses of neutron-deficient, 149Tm and 150Tm, combined with measurements of 149m,gEr (which were found to deviate from literature by $\sim$150 keV), provide the first experimental confirmation that 149Tm is the first proton-unbound nuclide in the Tm chain. Our measurements also enable the strength of the $N=82$ neutron shell gap to be determined at the Tm proton drip-line, providing evidence supporting its continued existence.
preprint: APS/123-QED
I Introduction
When isotopes become sufficiently neutron-deficient it eventually becomes energetically possible for one or more protons to spontaneously escape the nucleus. This transition within an isotopic chain to one or more protons being unbound is known as the proton drip-line, and its location is of fundamental interest. It can be experimentally determined through measurements of the one-proton separation energy, $S_{p}$, and a nucleus is said to lie beyond the proton drip-line when $S_{p}\leq 0$ (i.e. it is proton-unbound) [2, 3]. Studying the location of the proton drip-line provides a valuable benchmark for the various nuclear models which predict the properties of unstable nuclei.
In order for a proton to be emitted, it first must tunnel through the Coulomb barrier (i.e. a single-nucleon analogue to $\alpha$-decay). In contrast to alpha decay, it is however simpler to calculate the proton emission decay rate or spectroscopic factors as it doesn’t require the formation of a cluster of nucleons [2]. On the other hand, proton emission is significantly influenced by the centrifugal barrier and the proton emission half-life at a given proton number, $Z$, depends strongly on the underlying nuclear structure, the angular momentum carried away by the emitted proton [2][4][5], and the decay energy, $Q_{p}$ [5].
Furthermore, it has become evident that the nuclear shell structure can change towards the drip-line. Effects such as the weakening or disappearance of shells (quenching) at the classical magic numbers and the appearance of new magic numbers have been predicted and observed [6, 7]. Nuclear masses and binding energies provide indispensable information when studying the limits of nuclear existence. They allow calculations of the particle separation energies and can be used to make predictions regarding the underlying nuclear structure and nuclear properties such as half-life, angular momentum, or spectroscopic factors in the case of proton emission. Moreover, the proton drip-line around $Z=69$ is associated with an abrupt change in the deformation [8, 2, 9].
The mass values of nuclides across the $N=82$ shell closure, formed between the $h_{11/2}$ (or nearby $s_{1/2}$ or $d_{3/2}$) and the $f_{7/2}$ orbitals, have been intensively studied on the neutron-rich side of the nuclear chart, see e.g. [10, 11, 12, 13, 14, 15]. Corresponding investigations at the neutron-deficient $N=82$ shell closure up to the elements of holmium ($Z=67$) and erbium ($Z=68$) have been preformed by SHIPTRAP [16], ISOLTRAP [17], by the ESR using Schottky Mass Spectrometry [18] and recently in ytterbium ($Z=70$) at TITAN [19]. However masses of neutron-deficient thulium ($Z=69$) isotopes across $N=82$ remain elusive right where the one-proton drip-line is predicted to lie.
Here we report first time high precision mass measurements of neutron-deficient thulium isotopes, which close the gap across the proton drip-line and $N=82$ in the isotopic chain.
I.1 Background
Ground-state proton emission can occur for any nucleus beyond the proton drip-line, and was first observed in 1982 in 151Lu and 147Tm [20, 21]. Since then, at least 27 heavy ground-state proton emitters have been identified, alongside a number of isomeric states [8].
While 147Tm has long been known to be a proton emitter, and prior direct mass measurements of 147Tm and 148Tm have been able to constrain the location of the proton drip line to lie somewhere between $A=148$ and $A=151$ [16], the precise isotope at which Tm becomes proton-unbound remains uncertain due to the thus far unmeasured mass values of 149Tm and 150Tm. In cases where neither observation of the decay nor direct mass measurements can be used to establish the drip-line, various theoretical approaches are employed to predict its location but these models are known to struggle to predict $S_{p}$ far from stability [22]. Density functional theory and phenomenological descriptions of the nuclear binding energy have been utilized to predict the location of the proton drip-line. For instance, the Skyrme results of [23] can be used to interpolate the $S_{p}$ of odd-Z nuclei. However, for 149Tm, the choice of energy density functional – such as UNEDF1[24] or SV-min[25] – leads to competing predictions regarding its position relative to the drip-line. Furthermore, the newer BSkG3 model [26] places the drip-line even closer to stability.
Bayesian model averaging can be used to assign weights to models in an attempt to improve estimates of the binding energies of heavy nuclei [27]. A recent Bayesian statistical analysis was performed to weight the proton drip-line predictions provided by several different energy density functionals, as well as FRDM2012 and HFB-24 [28]. This work predicted the last proton-bound isotope of Tm to occur at $N=81$ ($A=150$), but this has never been experimentally verified.The precise location of the Tm drip-line therefore remains a significant hole in our experimental knowledge of the heavy, proton-deficient nuclei, and the present measurement aims to uncover the exact location of the ground-state proton drip-line of Tm ($Z=69$).
II Experimental Description
The isotopes under study were produced throughout two experimental campaigns at TRIUMF’s Isotope Separator and Accelerator (ISAC) facility[29] using the ISOL technique. A 480 MeV proton beam was impinged on a Ta target to produce rare isotopes via spallation reactions and Tm and Er isotopes were surface ionized in the target region. The continuous beam of neutron-deficient isotopes was purified using the ISAC high-resolution mass separator [30] (resolving power $\sim$2500) to select beams from $A=149$ to $A=157$ before preparation in the TITAN Radio-Frequency Quadrupole (RFQ) cooler-buncher trap[31] for precision mass measurements. The resulting bunches of ions were extracted and their kinetic energy matched for injection into TITAN’s MR-TOF-MS [32][33][34] to obtain a mass spectrum. In the MR-TOF-MS ions are again cooled and narrowly bunched before executing isochronous (energy independent) reflections between two electrostatic ion mirrors and being detected using a MagneTOFTM ion detector manufactured by ETP ion detect. All ions observed in the spectrum were delivered and measured in a 1+ charge state and any possible non-isobars present due to charge exchange were removed using a Mass Range Selector [35, 36]. An example spectrum of $A=149$ is given in Figure 1.
Measurements of the nuclei furthest from stability were made possible by employing so called mass selective re-trapping [37, 38] to selectively reduce the intensity of unwanted contaminant species [19]. This improved the signal-to-background ratio of the 149,150Tm measurements and allowed the MR-TOF-MS to accept a higher overall beam rate while keeping the total number of ions detected per mass measurement cycle below one.
During the initial experiment (experiment #1), isotopes were produced using a high-power tantalum target which is equipped with greater heat dissipation, whereas the second experiment (experiment #2) utilized a low-power tantalum target, for which an improved and more homogeneous thermal profile is believed to have enhanced the release of Tm and Er. In both experiments the proton beam from TRIUMF’s main driver cyclotron was impinged on the Ta production targets with beam currents ranging from approximately 25 to 50 $\mu$A.
The mass of 150Tm was measured alongside Yb masses [19] in experiment #1 which additionally provided anchor masses for alpha decay chains [39]. The resolving power of 270k in this experiment resulted in a proton separation energy of 150Tm which was very close to zero, and no definitive conclusion could be drawn as to whether this nucleus is proton-bound. The Tm data-set was greatly improved during experiment #2 by performing additional measurements extending measurements to 149Tm, and improving on the initial experimental uncertainties. Particularly, experiment #2 benefited from improvements that were made to the MR-TOF-MS system to achieve 400k resolving power and lower systematic uncertainty.
Figure 1: Top: The $A=149$ spectrum with major peaks labeled. 149Dy is used as both shape and mass calibrant here ($m$ indicates a long-lived isomer). Bottom: A fit of the 149g+mEr peak using the hG functional form and calibrated to the shape of the higher statistics 149Dy peak. Use of the hyperEMG model enabled the mass of the ground-state to be directly determined alongside the isomer within the mixed peak.
II.1 Analysis Procedure
The analysis procedure followed the description in [40]. At all mass units, we utilized a high abundance peak to correct for time-dependent drifts in the Time of Flight (TOF) spectrum using in-house developed data acquisition software[33]. Subsequently, masses were determined by fitting the peaks in each spectrum using a hyper-exponentially-modified Gaussian (hyperEMG) lineshape[41]. This procedure involves first fitting an isolated high-statistics peak with a multi-dimensional function to precisely determine the shape (i.e. functional form) of a single peak (the shape calibrant) and then using multiple copies of this shape to fit all peaks in the spectrum. Next, the masses of all peaks in the spectrum were calibrated against a single species with a well-established mass (the mass calibrant). Appropriate shape and mass calibrant peaks were chosen for each set of isobars measured to minimize the influence of overlapping isobars and nuclear isomers. In the present analysis, the peaks corresponding to 149Dy, 150Ho, 151Ho, 152Dy, 153Dy, 154Dy, 155Yb, 156Yb, and 157Yb were selected as shape calibrants in their respective isobaric mass spectra. The chosen mass calibrant at each mass number was cross-checked for consistency with the masses of several different isobars listed in the 2020 Atomic Mass Evaluation (AME2020) to confirm its identity, and are given in Table 1. However, multiple species contained long lived isomers which were not resolvable from their respective ground-state. Such unresolved contributions may shift the resulting mass values to higher mass as would be expected for a pure ground-state of the given isotope. When the excitation energy is known, we apply a correction and additional uncertainty following the guidelines laid out in the AME2020 appendix for the treatment of unresolved peaks. However, according to ENSDF ${}^{151,152,154}_{~{}82,~{}83,~{}85}$m respectively contain a $\sim 7$ s, a $5.2$ s, and a $3.3$ s isomer of unknown excitation energy which are likely to have influenced the measurement [42, 43, 44]. The mass values corresponding to the mixed peaks are included for completeness.
II.2 Results
In this work, mass measurements across $17$ neutron-deficient Tm and Er isotopes were obtained, with 149Tm, 150Tm representing first-time measurements. Mass excess values are defined as the sum of the ionic mass and the electron mass (since all ions were singly-charged) minus the mass number $A$.
The final mass values were obtained by adding statistical, fit and known systematic uncertainties in quadrature and are summarized in Table 1. Mass excess (ME) uncertainties include the uncertainty of the calibrant mass, while the Ion Mass Ratio uncertainties do not. The systematic uncertainty was dominated by non-ideal switching of voltages in the second ion mirror. For the initial measurement of 150Tm presented here a systematic uncertainty of $3\times 10^{-7}$ was included, whereas for all other measurements obtained during experiment #2 a systematic uncertainty of $1.5\times 10^{-7}$ was incorporated. This reduction was the result of improvements made to the tune of the MR-TOF-MS and the stability of the high voltage electrodes.
The crowded $A=153$ spectrum was the most challenging from which to obtain a high-statistics shape calibrant in the analysis. An isobaric species lies in each of the two tails of the 153Dy shape/mass calibrant, and the functional form needed to be manually selected in order to ignore excessive influence to the fit from the isobar in the lower mass tail. Furthermore, having a total rate of $\sim$3.8 ions per trapping cycle, with $\sim$2.7 per cycle in the 153Dy calibrant peak, necessitated the inclusion of an additional systematic uncertainty of $8\times 10^{-8}$ per ion per cycle to account for ion-ion interactions.
The species 154Tm, 152Tm, and 151Tm each contain a long-lived isomer for which the excitation energy is unmeasured. Since both the excitation energy and the ratio of ground-state to isomer entering the MR-TOF-MS are unknown, it is difficult to assign a definite mass for either ground-state or isomer, but the results are included for completeness.
153Tm has a long-lived isomer at 43.2$\pm$0.2 keV and 155Tm one at 41$\pm$6 keV. Both mass peaks are likely an unresolved admixture of isomer and ground-state. As per the standard procedure in Appendix B of AME 2020[45], the ground-state mass ($m_{gs}$) was obtained from the experimental mass ($m_{exp}$) and the known excitation energy ($\Delta E$) through the relation $m_{\textrm{ground-state}}=m_{\textrm{experiment}}-\Delta E/2$, and a corresponding additional uncertainty ($\delta m_{\textrm{admixture}}$) was included, where $\delta m_{\textrm{admixture}}=0.290(\Delta E)$.
The mass of 157Lu was likewise corrected for an unknown contribution from a 20.9$\pm$2.0 keV isomer (4.8 s) to obtain a mass 56$\pm$29 keV heavier than the AME value.
The mass spectra of 149,150Tm strongly benefited from mass selective retrapping improving the Tm-to-contaminant ratio. Once applied the peaks were well separated and could be fitted. No indication of the 670 keV, 5 ms 150Tm isomer was seen in the mass spectrum.
Overall we see good agreement between our new mass values and literature values from AME2020, particularly for 157Tm and 156Tm which are not known to have long-lived isomers or isobaric species that could impact our measurements. The two new Tm mass values now close the gap between a series of previously measured neutron deficient isotopes [16, 46, 47] and the neutron-rich side of the nuclear chart, with mass data now spanning from N=78 (147Tm) all the way to N=99 (168Tm). Alongside the measured Tm isotopes, 149-154Er were also identified in the mass spectra. The masses of 150-154Er all lie within 1.5 $\sigma$ of the presently accepted values in the AME2020 and any known isomers are $>2.5$ MeV above the ground state and thus do not interfere with the measurements.
149gEr and 149mEr were fit using two hyperEMG functions calibrated to the 149Dy peak (Fig. 1) to account for its known $11/2^{-}$ isomeric state, which was dominantly delivered to TITAN. Our 149gEr and 149mEr masses were found to be heavier by 2.9 $\sigma$ and 3.8 $\sigma$ respectively when compared to the literature value obtained via a measurement of 149mEr at the GSI ESR [18]. This results in a shift of the $S_{p}$ of 150Tm (see below).
Table 1: Mass measurements with their respective calibrants, ion mass ratios, and mass excess results compared to AME2020 values. The 150Tm mass from experiment #1 [19] is included at the bottom. The # indicates an extrapolated mass in the AME2020. Nuclides with a * have a long-lived isomer of unknown excitation energy, which may influence the measurement to an unknown degree. No correction was made and no additional uncertainty added and the mass excess of the admixture is reported (in italics). Nuclides with a contained an unresolved admixture of the ground-state and an isomer of a known excitation energy in the peak; for these a correction was made to the mass excess and the ion mass ratio (in bold) and an additional uncertainty added to reflect this (see section III). The 155Tm peak has an unresolved 41(6)keV isomer [48], 153Tm has a 43.2$\pm$0.2 keV isomer [49], and 157Lu has a 20.9$\pm$2.0 keV isomer [50]. As described in the text, these values were corrected using the procedure outlined in the 2020 Atomic Mass Evaluation (AME2020) [45] to obtain a mass for the ground-state and its uncertainty. A corresponding additional uncertainty of 13 keV for 153gTm and 12 keV for 155gTm is included. The high rate of 153Dy also necessitated the inclusion of an additional systematic uncertainty of $3\times 10^{-7}$.
[…]
157gLu † 157Yb 1.00004813(16) -46384 (26) -46440 (12) 56(29)
157Tm 157Yb 0.99996379(18) -58714 (28) -58709 (28) -4(40)
156Tm 156Yb 0.99997554(15) -56818 (24) -56834 (14) 17(28)
155gTm † 155Er 1.00003875(17) -56617 (26) -56626 (10) 9(28)
155Yb 155Er 1.00008157(15) -50437 (23) -50503 (17) 66(29)
154g+mTm * 154Dy 1.00011179(18) -54366 (27) -54427 (14) 61(30)
154Er 154Dy 1.00005458(16) -62569 (24) -62605 (5) 36(25)
153gTm † 153Dy 1.00010661(38) -53957 (54) -53973 (12) 16(55)
153Er 153Dy 1.00006064(34) -60517 (48) -60467 (9) -50(49)
152g+mTm * 152Dy 1.00013096(23) -51585 (33) -51720 (50) 136(60)
152Er 152Dy 1.00006789(18) -60510 (26) -60500 (9) -10(27)
151g+mTm * 151Dy 1.00012821(17) -50728 (24) -50772 (19) 44(31)
151Er 151Dy 1.00007472(15) -58248 (22) -58266 (16) 19(27)
150Tm 150Dy 1.00016386(23) -46427 (33) -46490 (200#) ##
150Er 150Dy 1.00008237(16) -57807 (23) -57831 (17) 25(29)
149Tm 149Dy 1.00017251(24) -43762 (34) -43940 (200#) ##
149gEr 149Dy 1.00010171(33) -53584 (47) -53742 (28) 158(55)
149mEr 149Dy 1.00010695(17) -52857 (25) -53000 (28) 143(38)
150Tm (Exp 1) 150Dy 1.00016407(34) -46397 (47) -46490 (200#) ##
III Discussion
As detailed below, our measurements of the Tm and Er mass chains establish 149Tm as the first proton-unbound Tm isotope. Furthermore, they support recent evidence for the persistence of the shell closure at $Z=70$ [19] with a corresponding measurement of the trend in $\Delta_{2N}$ at $Z=69$.
Last Proton-Bound Isotope Determined from 149m,149gEr and 150gTm
The Er isotopic chain was measured alongside Tm during experiment #2 from 154Er to 149Er.
During this second measurement, a discrepancy in the mass of 149gEr relative to the literature value was found which significantly affects the one-proton separation energy ($S_{p}$) of the 150Tm ground-state. Together, the newly-measured 150Tm mass and the AME2020 mass of 149gEr results in a $S_{p}$ that is consistent with zero ($1\sigma$), but using both updated masses yields a positive proton separation energy consistent with proton-bound 150Tm.
As with 151Yb for which the ISAC target was seen to preferentially deliver the spin 11/2 isomer to the ground-state, 149Er, having two fewer protons similarly appeared to prefer the excited state. We therefore fit two hyperEMG functions to the 149Er peak observed at TITAN to account for both the dominant isomeric state ($t_{1/2}=8.9\pm$0.2 s), as well as the ground-state ($t_{1/2}=4\pm$2 s) which shows up as an excess within the low-mass tail.
The ground-state mass of 149Er can be determined by two different methods using the present data. Firstly, it can be established from the measured isomer mass and known excitation energy as was previously done in an experiment using Schottky mass spectrometry [18], which reported an uncertainty of $\pm$28 keV. A second method is to directly fit the ground-state within the tail of the isomer’s peak by first fitting 149Dy to a hyperEMG function to establish the peak shape.
The isomeric excitation energy needed for the first method was previously measured via a two $\gamma$-ray internal transition to the ground-state proceeding through a 630.5 keV (M4) $\rightarrow$ 111.0 keV (M1) cascade[51], and was later confirmed in subsequent experiments which yielded 630.3 keV $\rightarrow$ 111.3 keV [52] and 630.5$\pm$2.6 keV $\rightarrow$ 111.3$\pm$1.1 keV by [53]. The 149mEr isomer is now established to lie 741.8$\pm$0.2 keV above the ground-state according to the NUBASE 2020 evaluation [54].
This excitation energy is consistent with our recent result in which the newly measured mass of 150Yb was used to support a theoretical explanation for the remarkably consistent first excitation energy (approximately 750 keV) in the isotonic chain containing 149Er [19]. This has been attributed to a clustering of single-proton energy levels for the mildly deformed nuclei at $N=81$ [19] resulting from the single hole in the $h_{11/2}$ neutron orbital common to 149Er and many of its isotones [51]. This provides increased confidence in the overall reliability of the established 149mEr excitation energy, from which our mass measurement of 149mEr yields a 149gEr mass excess of -53599$\pm$25 keV, a 143 keV increase from the AME2020 value when using the mass of the long-lived 741.69$\pm$0.23 keV isomer (9.6$\pm$0.6 s) and its presently accepted excitation energy [55] instead of directly measuring the ground-state.
However, the present dataset also enabled a direct measurement of the ground-state mass by utilizing the lower abundance, overlapping TOF peak of the ground-state and employing the hG procedure. Directly fitting the lower abundance 149gEr (as in Figure 1) with the hG function calibrated to the shape of the 149Dy peak yields a mass excess of -53584$\pm$47 keV, an increase in mass of 158$\pm$55 keV from the AME2020 value. The excitation energy of 149mEr can then be determined from the difference between the measured ground-state and isomer to be 726$\pm$53 keV.
While the newly determined mass of 149gEr and the isomer 149mEr deviate from the mass value in the AME 2020, they are consistent with observable trends in the binding energy of nuclides in this region. The cause of our discrepancy in our measurement of this single mass is unknown, but may warrant a third measurement of the 149g,mEr isotope for confirmation.
These new masses of the 149Er isomer and ground-state lead us to adopt a revised one-proton separation energy of the ground-state which directly impacts the location of the proton drip-line in Tm, since the proton separation energy of a nucleus depends on both the mass of the candidate proton emitter as well as on the mass of the daughter nucleus of the decay. Both the 149,150Tm masses as well as the revised 149Er value were thus needed for an unambiguous identification of the Tm drip-line, since the new Tm masses alone did not exclude the possibility that 150Tm could also be proton-unbound when using the 149Er value from the AME2020. These measurements confirm that 150Tm is indeed proton-bound, agreeing qualitatively with the Bayesian model averaging predictions of [28] that 149Tm is the first proton-unbound thulium isotope. Furthermore, the proton emission Q-value of 149Tm has now been determined and can be used as an input for decay rate calculations.
149Tm: The First p-Unbound Isotope
While 150Tm was shown to be proton-bound, our measurement of 149gTm establishes it as the first proton-unbound isotope in the thulium chain, having a Q-value of +428 keV when calculated using our Tm measurement and the masses of 148Er and 1H from the AME2020. Using $Q=M_{parent}-\Sigma M_{products}$ gives $Q_{p}=M_{{}^{149}Tm}-M_{{}^{148}Er}-M_{{}^{1}\textrm{H}}=428(35)\textrm{keV}$, a decay energy that is on the high side of the AME extrapolation of 250$\pm$201 keV. This value can be used in estimations of the proton emission half-life (e.g. [4, 5]).
[…]
Figure 2: The one-proton separation energy ($S_{p}$) of Tm isotopes calculated from the present measurements of masses in the Tm ($Z=69$) and Er ($Z=68$) isotopic chains is shown in blue and connected by solid lines (The 150Tm data point from experiment #1 is offset horizontally for clarity). This includes our new mass of 149Er at $N=81$, but masses from the 2020 AME [55] were used for 1H, 148Er ($N=80$), and 155Er ($N=87$) in calculating $S_{p}$ of Tm. Theory calculations of $S_{p}$ are shown for comparison. Note that the unresolved admixtures of isomeric states of unknown excitation energy in 151,152,154Tm ($N=82,83,85$) may artificially lower their separation energy.
Aside from BSkG3 [26], the models in Figure 2 systematically over-predict $S_{p}$. Thus the present measurements demonstrate that 149Tm is the first proton-unbound nucleus, rather than 148 or 150Tm.
Evolution of the $N=82$ neutron shell
In addition to determining the location of the drip-line, the new masses of 149/150Tm allow for a comparison to predictions of the empirical shell gap of the $N=82$ shell closure defined by $\Delta_{2N}=ME(N,Z+2)+ME(N,Z-2)-2ME(N,Z)$ (see Figure 3). The models shown include UNEDF0 [56], UNEDF1 [24], HFB21 [57], Duflo-Zuker [58], FRDM [59], SkP [60], SV-min[25], and ETFSI-Q [61]. In this region most models over-predict the shell gap of the N=82 isotone chain by 0.5-3 MeV with the exceptions of Duflo-Zuker under-predicting by nearly 1 MeV and HFB21 which lies within 500 keV of the isotones from $Z=67$ to $70$.
The present work measures the $\Delta_{2N}$ of Tm at the N=82 shell closure and shows that it is, in fact, the first isotone to depend on the binding energy of a proton-unbound nucleus (149Tm). These measurements indicate the persistence of the $N=82$ shell closure, which is consistent with the recent corresponding measurement of Yb [19]. Tm continues the trend in the shell gap of the more stable $N=82$ nuclei to have a more rapid change in binding energy at even-N nuclei.
Figure 3: Comparison of models and experimental data for the empirical shell gap ($\Delta_{2N}$) split into two figures for clarity. The mass data from [62] and [19] was added for $Z=49$ and $Z=90$ respectively. Density Functional Theory and “liquid drop” models are shown.
IV Conclusion and Outlook
The masses of neutron deficient Tm and Er isotopes obtained in this work identify the precise location of the proton drip-line in the Tm isotopic chain. The transition from positive to negative proton separation energy is observed between 149gTm and 150gTm. Our updated mass of 149Er directly impacts the proton-bound nature of 150Tm since it would be the daughter nucleus of the proton emission if this decay mode were possible. The well-understood composition of the mass spectra provides confidence in our identification of isotopes using the MR-TOF-MS. The masses of two previously unmeasured Tm isotopes were measured in this experiment. Additonally, an update to the previously measured mass of 149mEr and a first direct measurement of the ground-state 149gEr determined them to be $\sim$150 keV heavier than the current literature values, thereby establishing 150gTm as proton-bound.
Together, our masses of 149gTm, 150gTm, and 149gEr establish that 149gTm, having 20 fewer neutrons than the only stable isotope of Tm (169Tm, $N=100$), is the first proton-unbound ground-state in this isotopic chain. This agrees with the drip-line prediction obtained from a recent Bayesian analysis of various theoretical models [28].
We directly measured the mass of 149gEr by performing hyperEMG fits on the time-of-flight spectrum obtained with TITAN’s MR-TOF-MS. Higher precision can be achieved by indirectly determining the ground-state mass from the isomer and the previously measured excitation energy (recent observations in the trends in the excitation energies of the (11/2-) isomers [19] have added credibility to this approach). This indirect approach is how the current literature measurement was obtained.
Our measured mass values provide valuable benchmarks for the models used to predict the binding energies of yet-unmeasured nuclei, allowing them to be compared and weighted appropriately.This experimental result can now provide useful feedback to the various theoretical models that use the proton drip-line as a benchmark for their predictive capabilities. As next-generation radioactive ion beam facilities begin to come online, our update to the literature masses highlights the importance of measuring and publishing not only the masses of previously unmeasured isotopes, but also of confirming the mass values in literature which provide integral inputs to models of both physical processes and nuclear properties.
This measurement highlights the challenges that can arise when identifying and measuring masses of heavier exotic species due to the abundance of isomers and possible isobars. Mass values taken from the AME2020 should be used with caution when used for training data or weighting of models.
V Acknowledgements
TITAN is funded by the Natural Sciences and Engineering Research Council (NSERC) of Canada and through TRIUMF by the National Research Council (NRC) of Canada.
This work was further supported by the German Research Foundation (DFG), Grant No. 422761894; by the German Federal Ministry for Education and Research (BMBF), Grants No. 05P16RGFN1 and No. 05P19RGFN1; by the Hessian Ministry for Science and Art through the LOEWE Center HIC for FAIR; by the JLU and GSI Helmholtzzentrum für Schwerionenforschung under the JLU-GSI strategic Helmholtz partnership agreement.
For the purpose of open access, the authors have applied a creative commons attribution (CC BY) licence to any author accepted manuscript version arising.